# Subgroups Matter for Robust Bias Mitigation

### v2

@@ 2.3 Bias mitigation without subgroups
We restrict our main analysis to established mitigation methods which use subgroup annotations because (a), they often serve as the upper bound for mitigation methods (Zhou et al., 2021; Pezeshki et al., 2021; Bayasi et al., 2024), (b), many nominally label-free methods still require subgroup annotations for validation or hyperparameter tuning (Pezeshki et al., 2024), and (c) label-free methods frequently infer subgroups, making it still essential to understand their role.
@@ 3.1.3 Subgroup definition across these methods
In this work, we distinguish between two categories of bias mitigation methods: reweighting (gDRO and resampling), and model-based methods (DomainInd and CFair). This is because of differences in how subgroups are generally defined. While in model-based methods subgroups are defined solely based on attribute(s) $A$, reweighting methods can additionally define subgroups over $\mathcal{A}\times\mathcal{Y}$. This is because the latter methods perform best if the each subgroup contains both positive and negative samples. We expand on this distinction and its implications in Appendix A.2.
@@ 3.2 Problem setting
We frame the task according to the fairness paradigm described in Jones et al. (2025), whereby the objective is to generalise from a biased training distribution to an unbiased testing distribution. Dataset bias can take many forms (as discussed in Jones et al. (2024)), but here, we focus on bias arising from spurious correlations between certain attributes of the data and the class label $Y$ which disappear in the unbiased deployment setting. We define two binary attributes $A$ and $S$ whose information is encoded in $X$ in the form of latent features $X_{A}$ and $X_{S}$ respectively. Features relating to the true class $Y$ are represented as $X_{Y}$. In the unbiased test setting, attribute-related information is independent of $Y$, such that $\mathcal{P}(Y)=\mathcal{P}(Y\mid A)=\mathcal{P}(Y\mid S)$. However, in $\mathcal{P}_{train}$, mechanisms like data selection may lead to the violation of this independence. Causal graphs illustrating this scenario are shown in Figure 1.
@@ 5.2 Groupings used for mitigation strongly impact bias mitigation performance
We find that test set performance is highly dependent on the subgroups used for mitigation, with some subgroups boosting overall performance by more than 0.10, while others reduce it by up to 0.04 (Figures 2 and A7). We first note that the four datasets reveal similar patterns across groupings for reweighting-based bias mitigation methods and model-based methods, suggesting that there are universal trends relating to the bias setting which generalise outside of the specific dataset and task. For reweighting-based methods, subgroups based on $(A,Y)$, i.e., $(A,Y)$, $(A,Y)_{8}$, $(Y,S,A)$, $(A,Y)$ with small levels of noise, and SC/no-SC subgroups, enable the model to focus on the $(A,Y)$ pairs without the spurious correlation during training. Therefore the model learns to predict $Y$ independently of $A$, leading to better generalisation performance. Conversely, subgroups which do not take $(A,Y)$ information into account tend to result in worse performance than the baseline model (for instance $S$,$(S,Y)$,$(S,Y)_{8}$,$Y$, $A$, and Random subgroups), as they fail to guide the model away from relying on the spurious attribute $A$. For model-based methods, a similar pattern is evident (Figure B7); $A$ subgroups generally present the best performance (analogous to $(A,Y)$ in data-based methods). For instance, with $A$ subgroups in DomainInd, performance is increased by 0.07 in the CXP dataset. On the other hand, $S$ subgroups consistently harm performance, decreasing it by up to 0.14 for CFair in MNIST.
@@
For example, there is little difference between results for $(A,Y)$ vs $(A,Y)_{8}$ and $(S,Y)$ vs $(S,Y)_{8}$ (Figure 2). This is most likely because the bias mitigation methods tested here are designed to maintain stability if there are many subgroups. In practice, this suggests that dividing subgroups into finer subgroups is unlikely to harm performance, and may even be beneficial if one is unsure which specific attributes may be responsible for underperformance. This insight could be relevant in settings where there are multiple levels of granularity to choose from (such as continuous data e.g. age, or multi-level categories e.g. ethnicity). We validate this with the real coarse and granular subgroups in Civil_comments, and indeed find no significant difference in unbiased generalisation for resampling and gDRO, as shown in Figure 3.

### v3

@@ 2.3 Bias mitigation without subgroups
We restrict our main analysis to established mitigation methods which use subgroup annotations because (a), they often serve as the upper bound for mitigation methods (Zhou et al., 2021; Pezeshki et al., 2021; Bayasi et al., 2024), (b), many nominally label-free methods still require subgroup annotations for validation or hyperparameter tuning (Pezeshki et al., 2024), and (c), label-free methods frequently infer subgroups, making it still essential to understand their role.
@@ 3.1.3 Subgroup definition across these methods
In this work, we distinguish between two categories of bias mitigation methods: reweighting (gDRO and resampling), and model-based methods (DomainInd and CFair). This is because of differences in how subgroups are generally defined. While in model-based methods subgroups are defined solely based on attribute(s) $A$, reweighting methods can additionally define subgroups over $\mathcal{A}\times\mathcal{Y}$. This is because the former methods perform best if each subgroup contains both positive and negative samples. We expand on this distinction and its implications in Appendix A.2.
@@ 3.2 Problem setting
We frame the task according to the fairness paradigm described in Jones et al. (2025), whereby the objective is to generalise from a biased training distribution to an unbiased testing distribution. Dataset bias can take many forms (as discussed in Jones et al. (2024)), but here, we focus on bias arising from spurious correlations between certain attributes of the data and the class label $Y$ which disappear in the unbiased deployment setting. We define two binary attributes $A$ and $S$ whose information is encoded in $X$ in the form of latent features $X_{A}$ and $X_{S}$ respectively. Features relating to the true class $Y$ are represented as $X_{Y}$. In the unbiased test setting, attribute-related information is independent of $Y$, such that $P(Y)=P(Y\mid A)=P(Y\mid S)$. However, in $P_{train}$, mechanisms like data selection may lead to the violation of this independence. Causal graphs illustrating this scenario are shown in Figure 1.
@@ 5.2 Groupings used for mitigation strongly impact bias mitigation performance
We find that test set performance is highly dependent on the subgroups used for mitigation, with some subgroups boosting overall performance by more than 0.07, while others reduce it by up to 0.08 (Figures 2 and A7). We first note that the four datasets reveal similar patterns across groupings for reweighting-based bias mitigation methods and model-based methods, suggesting that there are universal trends relating to the bias setting which generalise outside of the specific dataset, task, and mitigation method. For reweighting-based methods, subgroups based on $(A,Y)$, i.e., $(A,Y)$, $(A,Y)_{8}$, $(Y,S,A)$, $(A,Y)$ with small levels of noise, and SC/no-SC subgroups, enable the model to focus on the $(A,Y)$ pairs without the spurious correlation during training. Therefore the model learns to predict $Y$ independently of $A$, leading to better generalisation performance. Conversely, subgroups which do not take $(A,Y)$ information into account tend to result in worse performance than the baseline model (for instance $S$,$(S,Y)$,$(S,Y)_{8}$,$Y$, $A$, and Random subgroups), as they fail to guide the model away from relying on the spurious attribute $A$. For model-based methods, a similar pattern is evident (Figure B7); $A$ subgroups generally present the best performance (analogous to $(A,Y)$ in data-based methods). For instance, with $A$ subgroups in DomainInd, performance is increased by 0.07 in the CXP dataset. On the other hand, $S$ subgroups consistently harm performance, decreasing it by up to 0.14 for CFair in MNIST.
@@
For example, there is little difference between results for $(A,Y)$ vs $(A,Y)_{8}$ and $(S,Y)$ vs $(S,Y)_{8}$ (Figure 2). This is most likely because the bias mitigation methods tested here are designed to maintain stability if there are many subgroups. In practice, this suggests that dividing subgroups into finer subgroups is unlikely to harm performance, and may even be beneficial if one is unsure which specific attributes may be responsible for underperformance. This insight could be relevant in settings where there are multiple levels of granularity to choose from (such as continuous data e.g., age, or multi-level categories e.g., ethnicity). We validate this with the real coarse and granular subgroups in Civil_comments, and indeed find no significant difference in unbiased generalisation for resampling and gDRO, as shown in Figure 3.